# Over-the-Air Fusion of Sparse Spatial Features for Integrated Sensing and Edge AI over Broadband Channels

### v1

@@ Abstract
The sixth-generation (6G) mobile networks are differentiated from 5G by two new usage scenarios – distributed sensing and edge artificial intelligence (AI). Their natural integration, termed integrated sensing and edge AI (ISEA), promised to create a platform for enabling environment perception to make intelligent decisions and take real-time actions, which empowers wide-ranging applications, e.g., autonomous driving and collaborative robotics. A basic operation in ISEA is for a fusion center to acquire and fuse features of spatial sensing data distributed at many edge devices (known as agents). Its implementation is confronted by a communication bottleneck due to multiple access by numerous agents over hostile wireless channels. To overcome the bottleneck, we propose a novel framework, called Spatial Over-the-Air Fusion (Spatial AirFusion), which exploits radio waveform superposition to aggregate spatially sparse features over the air and thereby enables simultaneous access. The technology, which targets environment perception, is more sophisticated than conventional Over-the-Air Computing (AirComp) as it supports simultaneous aggregation over multiple voxels, which partition the 3D sensing region, and across multiple subcarriers. The efficiency and robustness of Spatial AirFusion are derived from exploitation of both spatial feature sparsity and multiuser channel diversity to intelligently pair voxel-level aggregation tasks and subcarriers with the objective of maximizing the minimum receive signal-to-noise ratio among voxels subject to instantaneous power constraints. Optimally solving the resultant mixed-integer problem of Voxel-Carrier Pairing and Power Allocation (VoCa-PPA) is a focus of this work. The proposed approach hinges on two useful results: (1) deriving the optimal power allocation as a closed-form function of voxel-carrier pairing and (2) discovering a useful property of VoCa-PPA that allows dramatic dimensionality reduction of the solution space. Building on the two results, both a low-complexity greedy algorithm and an optimal tree-search based approach are designed for VoCa-PPA. The latter features a customised design of a compact search tree and a fast tree search enabled by intelligent node pruning and agent ordering. Extensive simulations using real datasets demonstrate that Spatial AirFusion achieves significant reduction in computation errors and improvement in sensing accuracy as opposed to conventional AirComp without awareness of spatial sparsity.
@@ I Introduction
The sixth-generation (6G) mobile network warrants two essential capabilities, sensing and artificial intelligence (AI) [1]. The first capability involves the integration of diversified sensing modalities such as camera, mmWave, and LiDAR sensors to collect information from sensory data. The second capability is envisioned to support AI model deployments in 6G edge networks, enabling the delivery of intelligent services. Integrating these two essentials for advanced 6G applications, ranging from high-precision perception to human-machine symbiosis, leads to an emerging paradigm called Integrated Sensing and Edge AI (ISEA) [2]. In such a system, an edge device equipped with sensors, termed an agent, in a distributed sensing system acquires sensory data from its surroundings and sends features extracted using its local perception model to the edge server (i.e., fusion center) for aggregation and then inference to support a downstream AI application[3, 4]. However, ISEA faces a communication bottleneck due to the aggregation of high-dimensional sensing features over resource-constrained wireless channels [5, 6]. One promising solution for overcoming the bottleneck is called Over-the-Air Computation (AirComp), which exploits waveform superposition in simultaneous access to realize over-the-air data aggregation[7, 8, 9, 10]. Based on AirComp, we develop a novel framework, termed Spatial Over-the-Air Fusion (Spatial AirFusion), for communication-efficient multi-sensor fusion in environment perception over a broadband channel. Its distinctive feature is to exploit spatial feature sparsity and channel frequency selectivity to intelligently map voxels, which divide the sensing region, to subcarriers for performing voxel-level AirComp tasks. Thereby, the sensing performance is improved while computation complexity reduced.
@@
Precise environment perception underpins a set of killer application scenarios of 6G, e.g., autonomous driving and collaborative robots. State-of-the-art perception models [11] leverage LiDAR, mmWave, and camera data to generate spatial feature vectors associated with certain locations in the physical world, as opposed to location-agnostic features in conventional classification and object detection. This type of feature is known as voxel features, where one voxel represents a spatial region in an evenly spaced 3D grid of the sensing range[12, 13]. To support low-latency and large-scale environment perception in 6G networks requires task-oriented air-interface design targeting ISEA. As a specific use case of edge inference, ISEA can be implemented on the well-known split inference architecture[14, 15, 16, 17]. In this architecture, a global inference model is split into a device and a server sub-model with the former used for local feature extraction and the latter for remote inference[16]. It can be generalized to distributed split inference (for which ISEA is a special case) by deploying models at multiple devices for local feature extraction (or inference) and performing local-feature (or label) aggregation at the server to attain a high inference accuracy[15, 18, 19]. Most recently, Researchers have made an attempt on designing AirComp techniques to realize different feature-aggregation functions, which include maximization, in an ISEA system based on an end-to-end sensing performance metric[3].
@@
AirComp in its own right is a fast-growing area[7]. The principle of AirComp is to exploit the superposition of signals simultaneously transmitted by multiple agents such that the desired aggregation functions, e.g., averaging, multiplication, and maximization, can be realized over the air[20, 21]. To materialize accurate functional computation via AirComp requires coping with channel fading and noise. For this purpose, a line of techniques has been designed to minimize AirComp errors including optimal power control[22], multiple-input-multiple-output (MIMO) beamforming[23] and interference management[24]. Broadband transmission is prevalent in modern high-rate mobile systems, which is assumed in the current system model. This motivates researchers to study broadband AirComp by addressing issues such as power allocation among subcarriers[19, 25], subcarrier truncation to avoid deep fading[8] and exploitation of channel frequency diversity [26]. The co-existing information-transfer users and AirComp devices participating in federated learning are also studied where the rate-maximizing subcarrier allocation for the former is designed subject to a guarantee on the learning performance of the latter [27]. Latest research focuses on the application of AirComp to increasing the communication efficiency of federated learning (FL), creating an area called over-the-air FL (AirFL)[8, 9, 10]. In this paradigm, AirComp realizes over-the-air aggregation of local gradients or models uploaded by devices, from which the result is used to update a global model at an edge server[8, 9]. While traditional AirComp techniques aim at computation error minimization, the design objective of AirFL techniques is to accelerate learning and account for the specific characteristics of transmitted data (i.e., local gradients/models). This results in a rich set of task-oriented wireless techniques such as power control based on gradient statistics[28], data- and channel-aware sensor scheduling[29], adaptive precoding[30], etc.
@@
Existing studies on AirComp as discussed above all assume single-stream data sources without considering data spatial distributions. Nevertheless, spatial feature variation is a key characteristic of environment perception as reflected in two aspects. On one hand, spatial feature sparsity is due to the fact that features are dense only in voxels containing detectable objects (e.g., vehicles and pedestrians) but not others. On the other hand, spatial feature distributions as observed by different agents are heterogeneous because of their non-identical fields of perception and view angles. Another aspect of heterogeneity is multiuser channel diversity. One key effect of spatial feature variation is the spatial variation of AirComp error as elaborated in the sequel. Let the task of spatial feature aggregation be divided into voxel-level sub-tasks. Due to the sparsity and heterogeneity of spatial feature distributions, the subset of agents participating in aggregation varies from voxel to voxel, resulting in differences in corresponding AirComp errors as they depend on the numbers of participating agents (see, e.g.,[31]) and qualities of the associated channels. The errors can be manipulated using a mechanism called Voxel-Carrier (VoCa) Pairing that maps voxels to subcarriers for executing their sub-tasks. Via this mechanism, a large number of degrees-of-freedom due to numerous voxels and subcarriers can be exploited to improve the performance of Spatial AirFusion. Furthermore, VoCa Pairing can be integrated with power allocation over subcarriers to obtain additional performance gain, giving rise to the problem of optimal VoCa Pairing and Power Allocation (VoCa-PPA).
@@
  • Experiments. The performance of Spatial AirFusion is evaluated by extensive experiments using both synthetic and real datasets (i.e., OPV2V[33]). The proposed framework is demonstrated to outperform traditional AirComp without awareness of spatial sparsity by a large margin, e.g., 10 dB gain in AirComp error suppression and significantly improved end-to-end inference accuracy.
@@ II-A Agent Perception Model
Each agent is equipped with a LiDAR or camera sensor that has specific perception ranges in depth, height, and width. The resultant three-dimensional perception space is partitioned into a regular grid with each cell referred to as a voxel. The numbers of partitions along the depth, height, and width directions are denoted as $V_{\sf d}$, $V_{\sf h}$, and $V_{\sf w}$, respectively. Then the total number of voxels is given as $V=V_{\sf d}V_{\sf h}V_{\sf w}$, which is assumed identical for all agents for ease of notation. Then the voxels for each agent can be indexed by $v=1,2,...,V$. As illustrated in Fig. 1, each agent utilizes its voxel-perception model to generate an $L$-dimensional feature vector for every voxel to capture the spatial object information contained within the voxel, termed voxel feature vector [12, 11]. For voxel $v$, its feature vector on agent $k$ is denoted as $\mathbf{f}_{k,v}\in\mathbb{R}^{L}$. It can be a zero vector (i.e., $\mathbf{f}_{k,v}=\mathbf{0}$) if agent $k$ detects no objects in voxel $v$, which may be a true or false negative result due to occlusion or limitations of the agent’s sensor.
@@ II-B Cooperative Sensing Model
The agents upload their voxel feature vectors, $\{\mathbf{f}_{k,v}\}_{1\leq k\leq K,1\leq v\leq V}$, to the fusion center over wireless links. Considering an arbitrary voxel, say voxel $v$, the result from fusing the associated vectors is denoted as $\mathbf{g}_{v}$. For two representative fusion functions, namely average-pooling and max-pooling [^1] The proposed method can accommodate other fusion functions, e.g., weighted-sum pooling and square-root pooling, via linearly scaling each vector or employing the generalized AirPooling pre- and post-processing functions [3]., the $\ell$-th element of $\mathbf{g}_{v}$ is given as
@@ II-C Communication Model
Spatial AirFusion wirelessly implements the above feature-fusion process over a broadband channel, which is modeled as follows. The total bandwidth $B$ is partitioned into $M$ subcarriers using orthogonal frequency division multiplexing (OFDM). Without loss of generality, it is assumed that $M\geq V$, as otherwise transmission for all $V$ voxels can be carried out over multiple channel coherence blocks. The channel follows block fading where a subcarrier remains constant within a channel coherence block. It is assumed that the channel coherence time is not shorter than the duration of $L$ symbol slots and that channel state information (CSI) is available at the receiver and transmitters. This allows a voxel feature vector to be uploaded within a single channel coherence block. Assuming symbol-level synchronization (see [8] for synchronization techniques), all agents simultaneously transmit their feature vectors on assigned subcarriers. The $\ell$-th symbol received by the fusion center on the $m$-th subcarrier, $y_{m}[\ell]$, is given by
@@ II-D Performance Metric
The presence of channel distortion in Spatial AirFusion results in AirComp error, defined as the mean square error between the over-the-air aggregated data and the ground-truth fusion result[7]. Under per-agent power constraints, AirComp error, known to be inversely proportional to the receive SNR, is dominated by the worst channel due to the required magnitude alignment of received signals via channel inversion[22]. In the sensing context, the end-to-end sensing accuracy, prone to distortion in the aggregated intermediate features, has been shown to improve with the receive SNR in [3]. In AirFusion, we denote the receive SNR for the sub-task of aggregating voxel $v$’s features as $\gamma_{v}$. It is determined by the coordinated SNR level for its assigned subcarrier, i.e., $\gamma_{v}=P_{\mathsf{rx},m(v)}$ if subcarrier $m(v)$ is assigned for voxel $v$. As the voxel $v$’s aggregated features encode information of sensed spatial objects, a low $\gamma_{v}$ implies high feature distortion and hence low object detection accuracy in voxel $v$. This can lead to catastrophic results in mission-critical tasks, such as missing a pedestrian or vehicle in autonomous driving. Hence, to guarantee satisfactory object detection accuracies over the entire sensing region, we propose to define the performance metric for Spatial AirFusion, denoted by $U$, as the minimum receive SNR across all voxels: $U=\min\limits_{v\in\{1,\ldots,V\}}\gamma_{v}$.

### v2

@@ Abstract
The sixth-generation (6G) mobile networks feature two new usage scenarios – distributed sensing and edge artificial intelligence (AI). Their natural integration, termed integrated sensing and edge AI (ISEA), promises to create a platform that enables intelligent environment perception for wide-ranging applications. A basic operation in ISEA is for a fusion center to acquire and fuse features of spatial sensing data distributed at many edge devices (known as agents), which is confronted by a communication bottleneck due to multiple access over hostile wireless channels. To address this issue, we propose a novel framework, called Spatial Over-the-Air Fusion (Spatial AirFusion), which exploits radio waveform superposition to aggregate spatially sparse features over the air and thereby enables simultaneous access. The framework supports simultaneous aggregation over multiple voxels, which partition the 3D sensing region, and across multiple subcarriers. It exploits both spatial feature sparsity with channel diversity to pair voxel-level aggregation tasks and subcarriers to maximize the minimum receive signal-to-noise ratio among voxels. Optimally solving the resultant mixed-integer problem of Voxel-Carrier Pairing and Power Allocation (VoCa-PPA) is a focus of this work. The proposed approach hinges on derivations of optimal power allocation as a closed-form function of voxel-carrier pairing and a useful property of VoCa-PPA that allows dramatic solution space reduction. Both a low-complexity greedy algorithm and an optimal tree-search algorithm are then designed for VoCa-PPA. The latter is accelerated with a customised compact search tree, node pruning and agent ordering. Extensive simulations using real datasets demonstrate that Spatial AirFusion significantly reduces computation errors and improves sensing accuracy compared with conventional over-the-air computation without awareness of spatial sparsity.
@@ I Introduction
The sixth-generation (6G) mobile network warrants two essential capabilities, sensing and artificial intelligence (AI) [1]. The first capability involves the integration of diversified sensing modalities such as camera, mmWave, and Light Detection and Ranging (LiDAR) sensors to collect information from sensory data. The second capability is envisioned to support AI model deployments in 6G edge networks, enabling the delivery of intelligent services. Integrating these two essentials for advanced 6G applications, ranging from high-precision perception to human-machine symbiosis, leads to an emerging paradigm called Integrated Sensing and Edge AI (ISEA) [2]. In such a system, an edge device equipped with sensors, termed an agent, in a distributed sensing system acquires sensory data from its surroundings and sends features extracted using its local perception model to the edge server (i.e., fusion center) for aggregation and then inference to support intelligent decisions and real-time actions for a downstream AI application[3, 4]. However, ISEA faces a communication bottleneck due to the aggregation of high-dimensional sensing features over resource-constrained wireless channels [5, 6]. One promising solution for overcoming the bottleneck is called Over-the-Air Computation (AirComp), which exploits waveform superposition in simultaneous access to realize over-the-air data aggregation[7, 8, 9, 10]. Based on AirComp, we develop a novel framework, termed Spatial Over-the-Air Fusion (Spatial AirFusion), for communication-efficient multi-sensor fusion in environment perception over a broadband channel. Its distinctive feature, differentiating it from conventional AirComp, is to exploit spatial feature sparsity and channel frequency selectivity to intelligently map voxels, which divide the sensing region, to subcarriers for performing voxel-level AirComp tasks. Thereby, the sensing performance is improved while computation complexity reduced.
@@
Precise environment perception underpins a set of killer application scenarios of 6G, e.g., autonomous driving and collaborative robots. State-of-the-art perception models [11] leverage LiDAR, mmWave, and camera data to generate spatial feature vectors associated with certain locations in the physical world, as opposed to location-agnostic features in conventional classification and object detection. This type of feature is known as voxel features, where one voxel represents a spatial region in an evenly spaced 3D grid of the sensing range[12, 13]. To support low-latency and large-scale environment perception in 6G networks requires task-oriented air-interface design targeting ISEA. As a specific use case of edge inference, ISEA can be implemented on the well-known split inference architecture[14, 15, 16, 17]. In this architecture, a global inference model is split into a device and a server sub-model with the former used for local feature extraction and the latter for remote inference[16]. It can be generalized to distributed split inference (for which ISEA is a special case) by deploying models at multiple devices for local feature extraction (or inference) and performing local-feature (or label) aggregation at the server to attain a high inference accuracy[15, 18, 19]. For communication-efficient feature aggregation, an AirComp-based general framework is proposed in [3] to realize different feature-aggregation functions, which include maximization, in an ISEA system based on an end-to-end sensing performance metric. As a simultaneous-access technology, AirComp promises to solve the scalability issues in ISEA, enabling low-latency device cooperation, which motivates us to further investigate AirComp for ISEA-based cooperative perception.
@@
AirComp in its own right is a fast-growing area[7]. The principle of AirComp is to exploit the superposition of signals simultaneously transmitted by multiple agents such that the desired aggregation functions, e.g., averaging, multiplication, and maximization, can be realized over the air[20, 21]. To materialize accurate functional computation via AirComp requires coping with channel fading and noise. For this purpose, a line of techniques has been designed to minimize AirComp errors including optimal power control[22], multiple-input-multiple-output (MIMO) beamforming[23] and interference management[24]. Broadband transmission is prevalent in modern high-rate mobile systems, which is assumed in the current system model. This motivates researchers to study broadband AirComp by addressing issues such as power allocation among subcarriers[19, 25], subcarrier truncation to avoid deep fading[8] and exploitation of channel frequency diversity [26]. The co-existing information-transfer users and AirComp devices participating in federated learning are also studied where the rate-maximizing subcarrier allocation for the former is designed subject to a guarantee on the learning performance of the latter [27]. Going beyond computation of generic aggregation functions, AirComp can be applied and tailored for specific AI computation tasks. This idea of task-oriented AirComp design originated in AirComp applications in federated learning (FL), which created an area called over-the-air FL (AirFL)[8, 9, 10]. In this paradigm, AirComp realizes over-the-air aggregation of local gradients or models uploaded by devices, from which the result is used to update a global model at an edge server[8, 9]. While traditional AirComp techniques aim at computation error minimization, the design objective of AirFL techniques is to accelerate learning and account for the specific characteristics of transmitted data (i.e., local gradients/models). This results in a rich set of task-oriented wireless techniques such as power control based on gradient statistics[28], data- and channel-aware sensor scheduling[29], adaptive precoding[30], etc.
@@
Existing studies on AirComp as discussed above all assume single-stream data sources without considering data spatial distributions. Nevertheless, spatial feature variation is a key characteristic of environment perception as reflected in two aspects. First, features are sparsely distributed in the voxel dimension. At the outputs of prevalent sensing models (e.g., VoxelNet[11] and PointPillars[31]), features for a given voxel are non-zero only if the voxel contains detectable objects in the physical world (e.g., vehicles and pedestrians). Consequently, only a small portion of all voxels are nonzero due to finite sensing ranges, view occlusion, and sparse scattering of objects in space. For example, both[11] and[31] report over 90% empty voxels. Second, spatial feature distributions as observed by different agents are heterogeneous because of their non-identical fields of perception and view angles. Another aspect of heterogeneity is multiuser and frequency diversities of wireless channels. One key effect of spatial feature variation is the spatial variation of AirComp error as elaborated in the sequel. Let the task of spatial feature aggregation be divided into voxel-level sub-tasks. Due to the sparsity and heterogeneity of spatial feature distributions, the subset of agents participating in aggregation varies from voxel to voxel. This results in different AirComp errors for different voxels as the errors depend on the numbers of participating agents (see, e.g.,[32]) and qualities of the associated channels. The errors can be manipulated using a mechanism called Voxel-Carrier (VoCa) Pairing that maps voxels to subcarriers for executing their sub-tasks. Via this mechanism, a large number of degrees-of-freedom due to numerous voxels and subcarriers can be exploited to improve the performance of Spatial AirFusion. Furthermore, VoCa Pairing can be integrated with power allocation over subcarriers to obtain additional performance gain, giving rise to the problem of optimal VoCa Pairing and Power Allocation (VoCa-PPA).
@@
  • Experiments. The performance of Spatial AirFusion is evaluated by extensive experiments using both synthetic and real datasets (i.e., OPV2V[34]). The benchmarking schemes include 1) naive AirComp which schedules all sensors for all voxels without sparsity awareness; 2) AirFusion-Vanilla which adopts the sparsity-aware framework but randomly pairs voxels with subcarriers; 3) digital air interface where devices transmit orthogonally. The proposed framework is demonstrated to outperform naive AirComp by a large margin, e.g., 10 dB gain in AirComp error suppression and significantly improved end-to-end inference accuracy. Compared with digital air interface, AirFusion achieves $5.74$ times reduction in communication latency with the same inference accuracy.
@@ II-A Agent Perception Model
Each agent is equipped with a LiDAR or camera sensor that has its own perception range and perspective. Prior to fusion, each agent calibrates timestamp differences and performs local perspective transformation to project its view onto the FC’s coordinates based on the relative position and speed using existing techniques such as coordinate offsets [35] and AVR[36]. We thus assume a shared three-dimensional coordinate for all sensors, which is by convention partitioned into a regular grid with each cell referred to as a voxel. The numbers of partitions along the depth, height, and width directions are denoted as $V_{\sf d}$, $V_{\sf h}$, and $V_{\sf w}$, respectively. Then the total number of voxels is given as $V=V_{\sf d}V_{\sf h}V_{\sf w}$, and the voxels are indexed by $v=1,2,...,V$. As illustrated in Fig. 1, each agent utilizes its voxel-perception model to generate an $L$-dimensional feature vector for every voxel to capture the spatial object information contained within the voxel, termed voxel feature vector [12, 11]. For voxel $v$, its feature vector on agent $k$ is denoted as $\mathbf{f}_{k,v}\in\mathbb{R}^{L}$. It can be a zero vector (i.e., $\mathbf{f}_{k,v}=\mathbf{0}$) if voxel $v$ is outside the perception range of agent or voxel $v$ is in the perception range but no objects are detected in voxel $v$ by agent $k$. Even in the latter case, the detection result may be false negative due to occlusion, noises or hardware imperfections of the agent’s sensor.
@@ II-B Cooperative Sensing Model
The agents upload their voxel feature vectors, $\{\mathbf{f}_{k,v}\}_{1\leq k\leq K,1\leq v\leq V}$, to the fusion center over wireless links. Considering an arbitrary voxel, say voxel $v$, the result from fusing the associated vectors is denoted as $\mathbf{g}_{v}$. For two representative fusion functions, namely average-pooling and max-pooling, the $\ell$-th element of $\mathbf{g}_{v}$ is given as
@@
Remark 1.
(Supported Aggregation Functions) In this paper, we have considered feature averaging or maximizing as the aggregation function, which can cover a majority of multi-sensor fusion schemes in cooperative perception for autonomous driving by up to a linear scaling at each sensor. For example, elementwise averaging/maximum is considered in F-Cooper and V2VNet, while weighted-sum fusion based on attention scores is adopted in Where2comm, BEVFormer and ActFormer (see the survey[37]). Our framework is extensible to many other types of fusion functions. If the function belongs to the family of nomographic functions, which includes, for example, square-root pooling and geometric mean, the extension is achieved by applying corresponding data pre- and post-processing at sensors and servers, respectively. For non-nomographic functions, approximations with nomographic functions can be designed by existing methods, e.g., [38].
@@ II-C Communication Model
Spatial AirFusion wirelessly implements the above feature-fusion process over a broadband channel, which is modeled as follows. The total bandwidth $B$ is partitioned into $M$ subcarriers using orthogonal frequency division multiplexing (OFDM). Without loss of generality, it is assumed that $M\geq V$, as otherwise transmission for all $V$ voxels can be carried out over multiple channel coherence blocks. The channel follows block fading where a subcarrier remains constant within a channel coherence block. It is assumed that the channel coherence time is not shorter than the duration of $L$ symbol slots and that channel state information (CSI) is available at the receiver and transmitters [^1] As a common assumption in existing broadband AirComp literature (see, e.g., [19, 27]), we assume reliable acquisition of CSI through downlink pilots and channel feedback via existing schemes such as frequency-domain interpolation[39] and limited feedback[40]. While dedicated pilot design and feedback schemes (see, e.g., [41]) can further mitigate the overhead, relevant discussions are beyond the scope of this paper. . This allows a voxel feature vector to be uploaded within a single channel coherence block. Assuming symbol-level synchronization (see [8] for synchronization techniques), all agents simultaneously transmit their feature vectors on assigned subcarriers. The $\ell$-th symbol received by the fusion center on the $m$-th subcarrier, $y_{m}[\ell]$, is given by
@@ II-D Performance Metric
The presence of channel distortion in Spatial AirFusion results in AirComp error, defined as the mean square error between the over-the-air aggregated data and the ground-truth fusion result[7]. Under per-agent power constraints, AirComp error, known to be inversely proportional to the receive SNR, is dominated by the worst channel due to the required magnitude alignment of received signals via channel inversion[22]. In the sensing context, the end-to-end sensing accuracy, prone to distortion in the aggregated intermediate features, has been shown to improve with the receive SNR in [3]. In AirFusion, we denote the receive SNR for the sub-task of aggregating voxel $v$’s features as $\gamma_{v}$, which controls the aggregation quality of ${\mathbf{g}}_{v}$. It is determined by the coordinated SNR level for its assigned subcarrier, i.e., $\gamma_{v}=P_{\mathsf{rx},m(v)}$ if subcarrier $m(v)$ is assigned for voxel $v$. The performance metric shall thus be a function of the received SNR levels $\{\gamma_{v}\}_{v=1}^{V}$. To determine its form for sensing performance maximization requires a closer look into the downstream region proposal network (RPN) [43] for object detection tasks. An important property in RPN inference is its locality. Specifically, RPN slides a small neural network over the aggregated feature map, which takes in a small spatial window of voxel features and outputs the detection results for the associated voxel. Mathematically, the detection result for voxel $v$, ${\mathbf{r}}_{v}$, is given by ${\mathbf{r}}_{v}=\text{RPN}\left(\{{\mathbf{g}}_{v}\}_{v\in\mathcal{N}(v)}\right)$, where $\mathcal{N}(v)$ denotes the spatially neighboring voxels of voxel $v$. Then the object detection results for the entire space is obtained by sliding over all $v\in V$. We can see that the detection results for a given voxel $v$ only rely on a small subset of voxel features with spatial locality. In mission-critical tasks where misdetection in a single voxel can be catastrophic (e.g., missing a pedestrian in autonomous driving), reliable detection is demanded for all voxels. Hence, it important to rein in the feature distortion by improving $\gamma_{v}$ for every voxel instead of simply controlling average distortion. We therefore propose to define the performance metric for Spatial AirFusion, denoted by $U$, as the minimum receive SNR across all voxels: $U=\min\limits_{v\in\{1,\ldots,V\}}\gamma_{v}$.